Grade 12th passGeneral Physics

What is the focal length of a double convex lens for which the radius of curvature of
each surface is 50 cm and the refractive index of glass is 1.50 ?

narayan
8 Years agoGrade 12th pass

1 Answer

Saurabh Koranglekar
6 Years ago

Use lens maker’s formula: (1/ f)=(n-1)[(1/R1)-(1/R2)] with n=1.50,R1=50cm and R2=-50 cm.
The result is :f=50cm.
5.0